OpenStudy (anonymous):

Solve the exponential equation. Express the solution set in terms of natural logarithms. 8^3x = 2.2

OpenStudy (anonymous):

is that 8 times 3x = 2.2?

OpenStudy (anonymous):

or 8 to the power of 3x= 2.2

OpenStudy (anonymous):

8 to the power of 3x=2.2

OpenStudy (anonymous):

\[8^{3x}-2.2\] \[3x=\frac{\ln(2.2)}{\ln(8)}\] byt the change of base formula so \[x=\frac{\ln(2.2)}{3\ln(8)}\]